# Leveraging automatic strategy discoveryto teach people how to select better projects

### v1

@@ 4 A new metareasoning algorithm for discovering decision strategies for human project selection
Previous metareasoning methods are unable to handle some of the intricacies of the project selection problem. Therefore, we developed a new strategy discovery method based on MGPO (Heindrich et al., 2023), which overcomes the limitations that prevent MGPO from being applicable to project selection. To reflect the commonalities and innovations, we call our new strategy discovery method MGPS (meta-greedy policy for project selection). Similar to MGPO, our method approximates the value of computation (VOC) (Russell and Wefald, 1991) by myopically estimating the immediate improvement in decision quality. Improving upon MGPO, MGPS calculates the myopic approximation to the VOC in a way that accounts for discrete criteria ratings, criteria scaling, and multiple sources of information with different costs and reliabilities.
@@
MGPS calculates a myopic approximation to the VOC of asking an expert about a specific criterion of a single project according to Algorithm 1. To account for discrete advisor outputs, Algorithm 1 iterates over the discrete set of possible ratings the expert might give and estimates the probability of each rating () and the belief update that would result from it . The probability of each rating is computed using the cumulative distribution function () of the belief state for the selected project’s criterion score (see Line 7). Here, the standard deviation of the likelihood function encodes the expert’s reliability, the prior () is the current belief about the project’s score on the evaluated criterion, and the weights convert the criteria into a common currency. The belief update that would result from the observation () is computed by applying the belief state update (see Line 13 and Equation 3). For the highest and lowest possible ratings, we instead calculate using the open interval (see Lines 9 and 11). The updated expected value of the belief state according to an observation is then calculated using Bayesian inference to integrate the new observation into the belief state (see Line 13).
The VOC calculation depends on the posterior belief states that would result from the different possible observations (), weighted by their probabilities. If the evaluated project currently has the highest expected reward (see Line 15), the VOC calculation expresses the probability of observing a value low enough that the second-best project becomes the most promising option (see Line 16). In the case where the evaluated project did not have the highest expected termination reward, the VOC calculation expresses the probability of observing a value high enough to make the evaluated project the most promising option (see Line 18). Finally, the cost of the requested expert is weighted using a free cost weight parameter and subtracted from the VOC estimate (see Line 20).
@@
The full meta-greedy policy consists of calculating the VOC for all possible meta-level actions and iteratively selecting the meta-level action with the highest VOC. If no action has a positive VOC, the termination action is chosen.
@@ 5.1 MGPS Tutor: An intelligent tutor for teaching people how to select better projects
Building on the choice tutor by (Heindrich et al., 2023), our tutor repeatedly asks the learner to choose from a pedagogically chosen set of decision operations (see Figure 1) that always includes the query that MGPS would have performed. Moreover, it leverages MGPS’s VOC calculation (Algorithm 1) to score the chosen query, and then provides binary feedback on its quality. If learners select a suboptimal expert, project, or criterion, they receive feedback indicating the correct choice and have to wait for a short time. The unpleasantness of having to wait serves as a penalty (Callaway et al., 2022). Otherwise, they receive positive reinforcement and the next choice is displayed. To receive positive reinforcement, the learner must select a query whose VOC is sufficiently close, as determined by a tolerance parameter , to the VOC of the optimal action. We set the tolerance to .
Our tutor teaches the strategy discovered by MGPS using a novel sophisticated training schedule, which fosters learning by incrementally increasing the complexity of the training task. This learning methodology is also known as shaping (Skinner, 1953), and has been successfully applied to teach decision strategies to humans (Heindrich et al., 2023). Our training schedule varies the numbers of projects, how many different expert assessments learners have to choose between, and the specific types of expert assessments offered as choices. In total, our tutor teaches the discovered project selection strategy using ten training trials. The first seven training trials use a smaller version of the project selection task with only two projects, while the last three trials use the full environment with five projects. The number of choices gradually increases throughout training from 1 in the first training trial to 9 in the last three training trials. The tutor varies the types of choices across trials. After an initial trial with only a single choice, the tutor offers choices that focus on different criteria within the same project for two trials. Then, the tutor offers choices that focus on different experts within the same project for two trials. The remaining trials combine both types of highlights while sometimes also featuring queries about different projects and also increasing the overall number of choices.
@@ 5.2 Evaluating the effectiveness of MGPS Tutor in a training experiment
We evaluated the participants’ performance using two measures: their RR-score and click agreement (Heindrich et al., 2023). RR-score’s are normalized by subtracting the average reward of a random baseline algorithm and dividing by the participant scores’ standard deviation. The random baseline algorithm is defined as the policy that chooses meta-level actions at random until the maximum number of decision operations is reached. Click agreement measures, how well participants learned to follow the near-optimal strategy discovered by our method. Specifically, we computed for each participant, which proportion of their information requests matched the action taken by the approximately resource-rational strategy discovered by MGPS.
@@
To understand what participants in the MGPS Tutor condition were taught, we inspected the strategy discovered by MGPS. We found that this strategy systematically asks the most reliable experts ( and ) to evaluate projects on the criterion with the highest decision weight (). It starts by asking expert to rate criterion for a randomly selected project. If that rating is below the maximum score, the same expert is immediately asked to rate the same criterion for a different project. If expert gives project the highest possible score on criterion (), the strategy requests a second opinion about it from expert . If expert also evaluates the criterion’s value as , the decision process ends and project is selected. If expert provides a rating below , expert is asked about a different project. This process is repeated until a project in which both experts and estimate is found, or the maximum number of decision operations () is reached. It is important to note that while this decision strategy is relatively easy to understand, describe, and execute, the task of discovering this specific strategy as more resource-rational than the vast number of possible other strategies is considerably more difficult.
@@ 6 Performance evaluation
The results reported in the previous section show that MPGS can discover project selection strategies that are more effective than the strategies people discover on their own. But how does its performance compare to other strategy discovery algorithms? To answer this question, we evaluated MGPS on a computational benchmark. We chose PO-UCT (Silver and Veness, 2010) for comparisons because it is an established baseline for metareasoning algorithms in partially observable environments (Heindrich et al., 2023) and the more specialized MGPO algorithm is not applicable to project selection. PO-UCT utilizes Monte Carlo tree search to simulate the effects of different actions, resulting in more accurate results with increased computation time, making it a useful baseline for MGPS’s computational efficiency and performance.
@@ 7 Conclusion
People’s decision-making is prone to systematic errors (Kahneman et al., 1982), and although people are happy to delegate some decisions, most CEOs are unlikely to let AI decide which projects their company should pursue. Moreover, people are reluctant to use normative project selection procedures because they tend to be more tedious (Schmidt and Freeland, 1992, Liu et al., 2017, Kornfeld and Kara, 2013). Motivated by people’s insistence on making their own decisions with simple heuristics, we leveraged AI to discover and teach prescriptive decision strategies adapted to human cognition that perform substantially better than people’s intuitive strategies but are nevertheless simple enough that people use them. To this end, we introduced a metareasoning method for leveraging AI to discover near-optimal decision strategies for human project selection. Modeling project selection through the lens of resource rationality allowed us to formulate a mathematically precise criterion for the quality of decision strategies for human project selection. We further developed an efficient automatic strategy discovery algorithm automatically discovers efficient strategies for human project selection. Our algorithm discovered decision strategies that are much more resource-rational than the strategies humans discovered on their own and the strategies discovered by a general-purpose algorithm for solving POMDPs (PO-UCT). Using the decision strategies discovered by our algorithm, we created a cognitive tutor that uses a shaping schedule and metacognitive feedback to teach the strategies to humans. In the training experiment, our cognitive tutor fostered significant improvements in participants’ resource rationality.
@@
A main limitation of our method is that it is unknown how precisely the environment parameters need to be estimated to construct the metareasoning task, which can prove especially problematic when there isn’t much data on past decisions. Future work could investigate and potentially address this issue by extending MGPS with a Bayesian inference approach to estimate uncertain parameters of the environment (Mehta et al., 2022). This extension would make MGPS robust to uncertainties in the underlying distributions of evaluation criteria or expert reliability, and make it possible to discover resource-rational strategies even when data to estimate environment parameters is limited.
Encouraged by the promising results from successfully teaching humans in our naturalistic model of project selection, we are excited about future work assessing the real-world impact of improving people’s decision-making by evaluating their decisions directly in the real world. Relevant real-world decisions could be a company deciding which research project to invest in, or a charity deciding which initiatives maximize impact and cost-effectiveness. Similarly, while our cognitive tutor proved to be effective in teaching humans critical aspects of MGPS’ discovered strategies, we are looking forward to future pedagogical work that improves and further evaluates the mechanisms by which the cognitive tutor teaches strategies. Additionally, we are also excited about potential future work that combines MGPS with AI-Interpret (Skirzyński et al., 2021) to automatically derive human-legible recommendations for how to make project selection decisions. Lastly, although MGPS performed very well on our benchmarks, MGPS’s myopic approximation could fail in more complicated scenarios with interdependent criteria. Such challenges could be addressed by solving meta-level MDPs with methods from deep reinforcement learning, for example, by utilizing AlphaZero (Silver et al., 2017).
Our results indicate that it is possible to use resource-rational analysis combined with automatic strategy discovery to improve human decision-making in a realistic project selection problem. As selecting projects is a common problem faced by both organizations and individuals, improving their decision strategies in this setting would have a direct positive impact. Specialized project selection tutors for specific types of project selection problems could be used to train current and future leaders within companies and organizations. Moreover, project-selection tutors could be integrated into MBA programs to teach future decision-makers efficient decision strategies as part of their education. We are optimistic that our general methodology is also applicable to other real-world problems, offering a promising pathway to teach people efficient strategies for making better decisions in other areas as well. Besides project selection problems, we believe our approach could be used to improve real-world decision-making in areas such as grant-making and policy decisions.

### v2

@@ 4 A new metareasoning algorithm for discovering decision strategies for human project selection
In this section, we present a new metareasoning algorithm based on MGPO (Heindrich et al., 2025). Previous metareasoning methods, including MGPO, are unable to handle some of the intricacies of the project selection problem. Therefore, we developed a new strategy discovery method that overcomes the limitations that prevent MGPO from being applicable to project selection. To reflect the commonalities and innovations, we call our new strategy discovery method MGPS (Meta-Greedy policy for Project Selection; see Algorithm 1). Like MGPO, MGPS approximates the value of computation (VOC) (Russell and Wefald, 1991) by myopically estimating the immediate improvement in decision quality. Improving upon MGPO, MGPS calculates the myopic approximation to the VOC in a way that accounts for discrete criteria ratings, criteria scaling, and multiple sources of information with different costs and reliabilities.
On a high level, MGPS functions by iterating over all possible meta-level actions, approximating their VOC, and iteratively selecting the computation with the highest VOC. When the selected operation requests additional information, the resulting information is used to update the belief state about the evaluated project according to Bayesian learning (Equations 2-3). When the selected operation is the termination operation, the policy selects the project with the highest expected utility according to the current belief state. The method by which MGPS determines when to perform which operation is described in Algorithm 1.
The VOC of requesting a rating is the difference between the expected utility of receiving the resulting information and its cost. MGPS calculates a myopic approximation of the VOC of asking an expert about a specific criterion of a single project according to Lines 1-21. Concretely, this means that the VOC is approximated by the improvement in expected termination reward that would result from receiving the requested piece of information and then selecting a project compared to selecting a project immediately without obtaining any additional information (Lines 22-26).
The value of requesting a rating depends on the possible posterior belief states that would result from its potential values () and their probabilities. Crucially, for a computation to have a positive myopic VOC, it must be possible for the resulting rating to change the decision-maker’s belief about which project is best. Therefore, the myopic VOC of requesting an expert evaluation chiefly depends on the probability that it will change which project appears to be best. MGPS calculates this probability based on the expected reward of the project that the expert is asked to evaluate () and the expected reward of the best alternative project (Lines 3-5). If the evaluated project currently has the highest expected reward (see Line 22), the VOC depends on the probability of observing a value that is so low that the currently second-best project would become the most promising option if that value were observed. The potential improvement from selecting a different project is calculated as the difference in expected rewards from the best alternative project and the current best project after integrating the new observation (see Line 23). If the evaluated project does not have the highest expected termination reward, the VOC depends on the probability of observing a value high enough to make the evaluated project the most promising option (see Line 26).
The cost of the requested expert is scaled using a free cost weight parameter and subtracted from the VOC estimate (see Line 29). The cost weight parameter is optimized using Bayesian optimization (Mockus, 2012) and enables MGPS to balance the number of selected computations in cases where a purely myopic approach might lead to suboptimal results.
@@
To account for the fact that the advisor ratings are discrete rather than continuous, MGPS iterates over the discrete set of ratings that the expert might give and estimates the probability of each rating (). Expert ratings are modeled using the cumulative distribution function () of a Normal distribution centered on the current belief state and using the combined standard deviation of the belief and the expert’s reliability (see Line 12). The probability of each discrete observation is then estimated by the probability of a sample from the assumed normal distribution falling into an interval around the observation. Concretely, MGPS uses the following intervals: for lowest possible value, for intermediate values, and for the highest possible value. For each observation, MGPS also calculates the belief update that would result from it (). The simulated update to the belief state is computed by applying the belief state update equation (Equation 2) to integrate the new observation with the prior belief according to the expert’s reliability (see Line 19).
@@
The full meta-greedy policy (Line 31) consists in calculating the VOC for all possible meta-level actions and iteratively performing the meta-level action with the highest VOC until no available expert advice has a positive VOC. At that point, the termination action is chosen and the project with the highest expected reward is selected.
@@ 5.1 MGPS Tutor: An intelligent tutor for teaching people how to select better projects
Building on the choice tutor by (Heindrich et al., 2025), our tutor repeatedly asks the learner to choose from a pedagogically chosen set of decision operations (see Figure 1) that always includes the query that MGPS would have performed. Moreover, it leverages MGPS’s VOC calculation (Algorithm 1) to score the chosen query, and then provides binary feedback on its quality. If learners select a suboptimal expert, project, or criterion, they receive feedback indicating the correct choice and have to wait for a short time. The unpleasantness of having to wait serves as a penalty (Callaway et al., 2022a). Otherwise, they receive positive reinforcement and the next choice is displayed. To receive positive reinforcement, the learner must select a query whose VOC is sufficiently close, as determined by a tolerance parameter , to the VOC of the optimal action. We set the tolerance to .
Our tutor teaches the strategy discovered by MGPS using a novel sophisticated training schedule, which fosters learning by incrementally increasing the complexity of the training task. This learning methodology is also known as shaping (Skinner, 1953), and has been successfully applied to teach decision strategies to humans (Heindrich et al., 2025). Our training schedule varies the numbers of projects, how many different expert assessments learners have to choose between, and the specific types of expert assessments offered as choices. In total, our tutor teaches the discovered project selection strategy using ten training trials. The first seven training trials use a smaller version of the project selection task with only two projects, while the last three trials use the full environment with five projects. The number of choices gradually increases throughout training from 1 in the first training trial to 9 in the last three training trials. The tutor varies the types of choices across trials. After an initial trial with only a single choice, the tutor offers choices that focus on different criteria within the same project for two trials. Then, the tutor offers choices that focus on different experts within the same project for two trials. The remaining trials combine both types of highlights while sometimes also featuring queries about different projects and also increasing the overall number of choices.
@@ 5.2 Evaluating the effectiveness of MGPS Tutor in a training experiment
We evaluated the participants’ performance using two measures: their RR-score and click agreement (Heindrich et al., 2025). RR-score’s are normalized by subtracting the average reward of a random baseline algorithm and dividing by the participant scores’ standard deviation. The random baseline algorithm is defined as the policy that chooses meta-level actions at random until the maximum number of decision operations is reached. Click agreement measures, how well participants learned to follow the near-optimal strategy discovered by our method. Specifically, we computed for each participant, which proportion of their information requests matched the action taken by the approximately resource-rational strategy discovered by MGPS.
@@
To understand what participants in the MGPS Tutor condition were taught, we inspected the strategy discovered by MGPS and visualized it as a flowchart in Figure 2. We found that this strategy systematically asks the most reliable experts ( and ) to evaluate projects on the criterion with the highest decision weight (). It starts by asking expert to rate criterion for a randomly selected project. If that rating is below the maximum score, the same expert is immediately asked to rate the same criterion for a different project. If expert gives project the highest possible score on criterion (), the strategy requests a second opinion about it from expert . If expert also evaluates the criterion’s value as , the decision process ends and project is selected. If expert provides a rating below , expert is asked about a different project. This process is repeated until a project in which both experts and estimate is found, or the maximum number of decision operations () is reached. It is important to note that while this decision strategy is relatively easy to understand, describe, and execute, the task of discovering this specific strategy as more resource-rational than the vast number of possible other strategies is considerably more difficult.
@@ 6 Performance evaluation
The results reported in the previous section show that MPGS can discover project selection strategies that are more effective than the strategies people discover on their own. But how does its performance compare to other strategy discovery algorithms? To answer this question, we evaluated MGPS on a computational benchmark. We chose PO-UCT (Silver and Veness, 2010) for comparisons because it is an established baseline for metareasoning algorithms in partially observable environments (Heindrich et al., 2025) and the more specialized MGPO algorithm is not applicable to project selection. PO-UCT utilizes Monte Carlo tree search to simulate the effects of different actions, resulting in more accurate results with increased computation time, making it a useful baseline for MGPS’s computational efficiency and performance.
@@ 7 Conclusion
This article presented a first proof-of-concept demonstration that AI can be applied to discover and teach cognitive strategies enabling people to make better decisions in the real world.
@@
To this end, we introduced a metareasoning method for leveraging AI to discover near-optimal decision strategies for human project selection. Modeling project selection through the lens of resource rationality allowed us to formulate a mathematically precise criterion for the quality of decision strategies for human project selection. We further developed an efficient automatic strategy discovery algorithm that automatically discovers efficient strategies for human project selection. Our algorithm discovered decision strategies that are much more resource-rational than the strategies humans discovered on their own and the strategies discovered by a general-purpose algorithm for solving POMDPs (PO-UCT). Using the decision strategies discovered by our algorithm, we created a cognitive tutor that uses a shaping schedule and metacognitive feedback to teach the strategies to humans. In the training experiment, our cognitive tutor fostered significant improvements in participants’ resource rationality.
Prior work has extended strategy discovery to partially observable environments, a critical advance for utilizing strategy discovery to improve human decision-making in the real world (Heindrich et al., 2025). While the intelligent tutor in prior work proved effective at modeling partially observable problems and improving human decision-making, it shared a key limitation with previous work in the field: it was evaluated on abstract and simplified toy problems. In this article, we demonstrated a viable pathway to bringing strategy discovery into the real world: using real-world data to construct metalevel MDPs of real problems, solving the metalevel MDP with strategy discovery methods, and teaching the discovered strategies to humans with intelligent tutors.
This approach might make it possible to extend the curricula of intelligent tutoring systems beyond the procedural skills that instructional designers are readily able to articulate. Our demonstration that this approach is applicable to a decision-making skill commonly taught in MBA programs could become a milestone on the path to integrating automatic strategy discovery into educational applications of intelligent tutoring systems.
@@
A main limitation of our method is that it is unknown how precisely the environment parameters need to be estimated to construct the metareasoning task. With a highly accurate environment model, it is more likely that people will be able to directly apply the learned strategies to analogous situations in the real world, while a more abstract representation of the real-world task would require significant transfer of the learned strategy. While limited evidence for strategy transfer exists (Callaway et al., 2022b), we believe modeling real-world problems as accurately as possible is the most likely path towards real-world impact. This can prove especially problematic when there isn’t much data on past decisions. Future work could investigate this issue with a sensitivity analysis and potentially address it by extending MGPS with a Bayesian inference approach to estimate uncertain parameters of the environment (Mehta et al., 2022). This extension would make MGPS robust to uncertainties in the underlying distributions of evaluation criteria or expert reliability, and make it possible to discover resource-rational strategies even when data to estimate environment parameters is limited.
Another simplification of the environment model is that all observations are generated by retrieving a noisy sample of the true reward distribution, which is likely overly optimistic of real-world situations. For example, experts in the project evaluation task might have biases that lead to systematic overvaluation of specific criteria. This limitation could be addressed by integrating a model of people’s biases, such as utility weighted sampling (Lieder et al., 2015).
Our method, MGPS, is limited by the assumption that each computation is independent from all other computations. Real-world situations might involve more complex, interdependent planning steps. While it is possible to integrate this into the metalevel-MDP, MGPS’s myopic approximation could fail in more complicated scenarios with interdependent criteria. Such challenges could be addressed by solving meta-level MDPs with methods from deep reinforcement learning, for example, by utilizing AlphaZero (Silver et al., 2017).
Similarly, while our cognitive tutor proved to be effective in teaching humans critical aspects of MGPS’ discovered strategies, we are looking forward to future pedagogical work that improves and further evaluates the mechanisms by which the cognitive tutor teaches strategies. An especially exciting direction could be to expand the tutor with automatically generated natural language descriptions for how to make project selection decisions, such as using AI-Interpret (Skirzyński et al., 2021) or large language models.
While our implementation of project selection demonstrated the viability of improving human decision-making in naturalistic tasks, we did not evaluate our method directly in the real world. The most critical direction of future work is therefore to assess the methodology in a real-world field test. We believe our method to be generally applicable to a wide range of real-world decision-making problems that follow the general structure of being able to improve the quality of a decision (i.e. being able to choose a better project after learning additional information) at the cost of expending additional planning effort (i.e. the cost of requesting an expert opinion). Building upon our project selection model, relevant real-world decisions could be a company deciding which research project to invest in or a charity deciding which initiatives maximize impact and cost-effectiveness.
Our results indicate that it is possible to use resource-rational analysis combined with automatic strategy discovery to improve human decision-making in realistic scenarios. While our main goal was to provide a proof-of-concept demonstration for the general approach, project selection is an important practical skill that is commonly taught in MBA programs. Future work could leverage automatic strategy discovery methods to design of intelligent tutoring systems for training executive and managerial decision making. The resulting educational technologies could be useful for equipping future leaders with more rational, yet practical, project selection strategies. More generally, we view our results as demonstrating the feasibility of (1) modeling good real-world decision-making within the rational metreasoning framework, (2) discovering efficient cognitive strategies for real-world problems, and (3) improving human decision-making by teaching them the discovered strategies. We are optimistic that this general approach can ultimately be applied to teaching cognitive skills, including mathematical problem-solving, decision-making, and computational thinking, in educational settings ranging from high school to executive MBA programs.